# On Experiments

### v1

@@ Abstract
The scientific process is a means for turning the results of experiments into knowledge about the world in which we live. Much research effort has been directed toward automating this process. To do this, one needs to formulate the scientific process in a precise mathematical language. This paper outlines one such language. What is presented here is hardly new. The material leans much on great thinkers of times past [1, 2, 3, 4, 5] as well as more modern contributions [6, 7, 8, 9, 10]. The novel contributions of this paper are:
@@
    A new, general data processing inequality.
@@
1 Introduction
The notion of experiment is a key component of the scientific process. Intuitively, experiments link the values of some unknown property of the world to the outcomes of the experiment. Performing an experiment provides information. Here we construct a general language for defining experiments, the information contained therein and means to qualitatively and quantitatively compare experiments. Information will be seen to be governed by transformation, the information present explained by those transformations that render it untouched. Mathematically, the study of transformations is a subset of category theory. Therefore we make allusions to this theory where appropriate, showing where some constructions in statistics can be seen as general constructions in category theory.
These ideas are hardly new, and we lean heavily on great thinkers of times passed. We do argue that the presentation of the ideas is novel, modern and concise. Of course we also offer some new results. The novel contributions of this paper are:
  1. 1.
    A new, general data processing inequality.
  2. 2.
    A bias variance decomposition for canonical losses.
  3. 3.
  4. 4.
    Means to calculate deficiency via linear programming.
While our ultimate goal is the comparison of experiments, we also provide results on the representation of loss functions, section 4. The results in this section have appeared previously as appendix C of [11]. We include them here for completeness.
@@ 2 The General Decision Problem
We phrase the general decision problem as a two player game between nature and a decision maker. Let $\Theta$ be a set of possible values of some unknown, and $A$ the set of actions available to the decision maker. The consequence of an action is measured by a loss function $L:\Theta\times A\rightarrow\mathbb{R}$. A negative loss represents a gain to the decision maker. The norm of a loss function is given by its largest possible consequence (positive or negative), $\left\lVert L\right\rVert_{\infty}=\max_{\theta,a}\lvert L(\theta,a)\rvert$.
Unknown to the decision maker is the exact value of $\theta$. To discover this, the decision maker is guided by experiments. Experiments loosely are a special kind of relationship between the set of unknowns and the outcome of the experiment. Due to limitations in the type/number of tests available to the decision maker, as well as potential errors in the measure apparatus, the exact identity of $\theta$ can remain a mystery in spite of observed data. We assume that the decision maker represents their uncertainty in $\theta$ via probability.
@@ 3 Probability
The loss function itself does not tell us how to act, rather it provides a loss profile, a function $L_{a}:\Theta\rightarrow\mathbb{R}$, for each action $a\in A$. This loss profile summarizes the loss incurred for each value of the unknown for a fixed action. Our knowledge of $\theta$ allows us to place a loss on each of these actions, via an expectation,
@@
which in general can be any function from loss profiles to the reals. Ultimately, we choose the action we think has the lowest expectation. Exactly how one maps loss profiles to their expectations is up to the individual decision maker. Here we place the restriction that the expectation be given by a probability distribution.
@@
Point one states that probability distrbutions are normalized, point two that they are positive. At times we use infix notation, with $\mathbb{E}_{P}f=\left<P,f\right$>. Define the set of all distributions on a set $X$ to be $\mathbb{P}(X)$. For any $x\in X$ define the point mass distribution $\delta_{x}$, with $\mathbb{E}_{\delta_{x}}f=f(x)$ for all functions $f$.
@@ 3.1 Measure Theoretic Technicalities
The language as developed is fine for dealing with finite sets, where elements of the dual are merely finite dimension vectors. When dealing with infinite sets, the issue of how to represent elements of $\left(\mathbb{R}^{X}\right)^{*}$ becomes more troublesome. The usual route is to only consider loss profiles that are measurable. Rather than all functions $\mathbb{R}^{X}$ we work with a sub algebra of measurable functions which we will call $\mathrm{Meas}(X,\mathbb{R})$. A probability distribution is than a positive, normalized, linear function $\mathrm{Meas}(X,\mathbb{R})\rightarrow\mathbb{R}$. Restricting oneself in this way affords the representation of probability distributions via the Lebesgue integral. Of course one could take our work as pertaining to $\mathrm{Meas}(X,\mathbb{R})$ rather than all of $\mathbb{R}^{X}$, without any real loss in the results. Our primary interest is when:
The Basic Sets of Study are Finite
@@
At times, the natural set of actions will not be a finite set, but rather a well behaved subset of $\mathbb{R}^{n}$.
@@ 4 Optimal Decisions
Returning to our decision maker, we assume that they represent their uncertainty in $\Theta$, via a probability distribution $P\in\mathbb{P}(\Theta)$, and that they choose their actions via minimizing the expected loss,
$\operatorname*{arg\,min}_{a\in A}\mathbb{E}_{P}L_{a}.$
@@ 4.1 Representing Loss Functions
Here we develop general representations of loss functions $L:\Theta\times A\rightarrow\mathbb{R}$. We assume that the sets $\Theta$ and $A$ are finite.
Even if one starts with a finite set of actions, in many statistical problems it is natural to consider distributions over unknowns $\mathbb{P}(\Theta)$.
@@
A proper loss takes a prediction $Q\in\mathbb{P}(\Theta)$, and then penalizes the decision maker according to how much weight their prediction assigned to the unknown $\theta$. Intuitively properness ensures that if the decision maker knows $P$, then they minimize their expected loss by reporting $P$. Proper losses constitute a well studied class of loss functions, that provide suitable surrogates for decision problems [12, 9, 13, 14, 10, 15, 10, 16].
Furthermore, we show how to render any proper loss convex through a canonical re-parametrization. This allows the use of tools from convex analysis [17, 18] to aid in calculating optimal actions.
@@ 4.2 Entropy from Loss
Rather than working with probability distributions, we take the route of [19] and work with un-normalized distributions. Denote the set of all unnormalized distributions on $\Theta$ by $\mathbb{P}^{+}(\Theta)$. For any loss function $L$, define the entropy ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$,
@@
A function $f:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ is 1-homogeneous if for all $x\in\mathbb{P}^{+}(\Theta)$ and for all $\lambda>0$,
$f(\lambda x)=\lambda f(x).$
@@ 4.3 Loss from Entropy
All loss functions give rise to an entropy. Conversely, the entropy encodes much information of its associated loss through its super-gradients, which include all the Bayes actions for the underlying loss.
For any distribution $P$, define the Bayes actions for $P$ as the set of minimizers,
@@
Denote the set of all super-gradients at a point $x$ by $\partial f(x)$, and the set of all super-gradients by $\partial f=\cup_{x}\partial f(x)$. For differentiable concave functions, super-gradients are the same as regular gradients [18]. 1-homogeneous functions afford a very simple representation via their super-gradients.
@@
Let
$f:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave 1-homogeneous function. Then for all $x$ and for all $v\in\partial f(x)$,
@@
which follows directly from the definition of a super-gradient at $x$ and the 1-homogeneity of $f$. Re-arranging yields, $\frac{1}{2}(f(x)-\left<x,v\right>)\geq 0$, which implies $f(x)\geq\left<x,v\right$>. Similarly,
@@
which follows directly from the definition of a super-gradient at $\lambda x$ and the 1-homogeneity of $f$. Re-arranging yields, $\frac{1}{2}(f(x)-\left<x,v\right>)\leq 0$, yielding $f(x)\leq\left<x,v\right$>. Combing the results gives $f(x)=\left<x,v\right$>.
To prove the second claim, we have for all $y$ and $\lambda>0$,
@@
where the first line is by definition, and the second is by 1-homogeneity. As $y$ is arbitrary, the claim is proved.
This theorem provides a corollary, that shows the super-gradients of a 1-homogeneous function have a property similar to properness.
@@
We now show that the partial loss of a Bayes action is a super-gradient of $\underline{L}$}.
@@
Let ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave, 1-homogeneous function. Then its canonical loss, ${\mathcal{L}}:\Theta\times\partial{\underline{L}}\rightarrow\mathbb{R}$ is given by, ${\mathcal{L}}(\theta,\zeta)=\zeta(\theta)$.
As will be shown, canonical losses can always be convexified. Furthermore, they maintain all of the properties of $L$ needed for assessing the quality of decisions.
4.3.1 The Bayes Super Prediction Set
The process of canonising a loss, i.e. going from,
@@ 4.3.1 The Bayes Super Prediction Set
can create extra partial losses/actions that were not originally available to the decision maker. However, they gain no benefit from these extra actions. From any entropy define the Bayes super prediction set,
@@
The Bayes super prediction set is precisely those partial losses that the decision maker need not use over the actions available to them, no matter the distribution $P$. The super prediction set is convex. Furthermore, the Bayes actions for $\mathcal{L}$} are the lower boundary of the super prediction set.
@@
Let ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave, 1-homogeneous function. Then $\zeta\in\partial{\underline{L}}$ if and only if,
@@
The proof is a straightforward application of 1-homogeneity and super-gradients.
@@
Let ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave, 1-homogeneous function and let ${\nabla{\underline{L}}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}^{\Theta}$ be a super-gradient function, ${\nabla{\underline{L}}}(\mu)\in\partial{\underline{L}}(\mu),\ \forall\mu$. Then,
@@
is a proper loss. Furthermore if $\underline{L}$} is strictly concave then $L$ is strictly proper.
@@ 4.4 Convexification of Losses in Canonical Form
The preceding shows how to construct losses, we begin with a concave 1-homogeneous function and take super-gradients. Focus now turns to their convexification. Once convexified, the decision maker gains access to the large and ever growing literature on the minimization of convex functions to aid in the calculation of optimal actions. We closely follow [10], with a focus on canonical losses. This streamlines the development. For example, for some proper losses lemma 12 fails to hold. Furthermore, our result on convexification of canonical losses (theorem 14), is to the best of our knowledge novel.
Recall $\bm{1}_{\Theta}\in\mathbb{R}^{\Theta}$ is the function that always returns $1$, and define $\bm{1}_{\Theta}^{\bot}$ to be its orthogonal complement in $\mathbb{R}^{\Theta}$, i.e. the functions $v\in\mathbb{R}^{\Theta}$ with,
$\left<\bm{1}_{\Theta},v\right>=\sum_{z\in\Theta}v(z)=0.$
@@
Proof  Fix $v$ and suppose there exists $\gamma_{1}$ and $\gamma_{2}$ with $\gamma_{1}<\gamma_{2}$ and $\gamma_{1}\bm{1}_{\Theta}+v,\gamma_{2}\bm{1}_{\Theta}+v\in\partial{\underline{L}}$. By assumption, $\gamma_{2}\bm{1}_{\Theta}+v$ is Bayes for some distribution $P$. But,
@@
Thus we lose nothing by working with projections of losses onto $\bm{1}_{\Theta}^{\bot}$. Define,
@@
Proof  To show $\hat{\Gamma}_{{\underline{L}}}$ is convex, we are required to show that for all $\zeta_{1},\zeta_{2}\in\partial{\underline{L}}$ and all $\lambda\in[0,1]$ there is a constant $\gamma$ such that,
@@
with equality for $P^{*}$. Therefore by lemma 9, $\lambda\zeta_{1}+(1-\lambda)\zeta_{2}-\gamma^{*}\bm{1}_{\Theta}\in\partial{\underline{L}}$.
@@
Let ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave, 1-homogeneous function. Then its canonical loss $\mathcal{L}$} can be represented as ${\mathcal{L}}:\Theta\times C\rightarrow\mathbb{R}$, with $C\subseteq\bm{1}_{\Theta}^{\bot}$ a convex set and,
${\mathcal{L}}(\theta,v)=-\left<\delta_{\theta},v\right>+\Psi(v),$

### v2

@@ Abstract
The scientific process is a means to turn the results of experiments into knowledge about the world in which we live. Much research effort has been directed toward automating this process. To do this, one needs to formulate the scientific process in a precise mathematical language. This paper outlines one such language. What is presented here is hardly new. The material is based on great thinkers from times past [19, 20, 2, 7, 12] as well as more modern contributions [13, 17, 16, 10, 6]. The novel contributions of this paper are:
@@
    A new general data processing inequality.
  2. 2.
    A bias variance decomposition for canonical losses.
  3. 3.
  4. 4.
    Means of calculating deficiency through linear programming.
1 Introduction
The notion of an experiment is a key component of the scientific process. Experiments link the values of some unknown property of the world to the outcomes of the experiment. Performing an experiment provides information. We construct a general language for defining experiments, the information contained therein, and means to qualitatively and quantitatively compare experiments. Information will be seen to be governed by transformation, the information present explained by those transformations that render it untouched. Mathematically, the study of transformations is a subset of category theory. Therefore, we make allusions to this theory where appropriate, showing where some constructions in statistics can be seen as general constructions in category theory.
These ideas are hardly new, and we lean heavily on great thinkers of the past. We argue that presentation of the ideas is novel, modern, and concise. Of course, we also offer some new results. The novel contributions of this paper are:
  1. 1.
    A new general data-processing inequality.
@@ 1 Introduction
While our ultimate goal is the comparison of experiments, in section 4 we also provide results on the representation of loss functions. The results in this section have appeared previously, as Appendix C of [18]. We include them here for completeness.
@@ 2 The General Decision Problem
We phrase the general decision problem as a two-player game between nature and a decision maker. Let $\Theta$ be a set of possible values of some unknown, and $A$ the set of actions available to the decision maker. The consequence of an action is measured by a loss function $L:\Theta\times A\rightarrow\mathbb{R}$. A negative loss represents a gain to the decision maker. The norm of a loss function is given by its largest possible consequence (positive or negative), $\left\lVert L\right\rVert_{\infty}=\max_{\theta,a}\lvert L(\theta,a)\rvert$.
Unknown to the decision maker is the exact value of $\theta$. To discover this, the decision maker is guided by experiments. An experiment is a special kind of relationship between the set of unknowns and the outcome of the experiment. Due to limitations in the type/number of tests available to the decision maker, as well as potential errors in the measurement apparatus, the exact identity of $\theta$ can remain a mystery despite observed data. We assume that the decision maker represents their uncertainty in $\theta$ through probability.
@@ 3 Probability
The loss function itself does not tell us how to act; it provides a loss profile, a function $L_{a}:\Theta\rightarrow\mathbb{R}$, for each action $a\in A$. This loss profile summarizes the loss incurred for each unknown value for a fixed action. Our knowledge of $\theta$ allows us to place a value on each of these actions, through an expectation,
@@
which generally can be any function from loss profiles to reals. Ultimately, the decision maker chooses the action that they think has the lowest expectation. Exactly how one maps loss profiles to their expectations is up to the individual. Here we place the restriction that the expectation should be given by a probability distribution.
@@
Point one states that probability distributions are normalized, point two that they are positive. At times we use infix notation, with $\mathbb{E}_{P}f=\left<P,f\right$>. Define the set of all distributions on a set $X$ to be $\mathbb{P}(X)$. For any $x\in X$ define the point mass distribution $\delta_{x}$, with $\mathbb{E}_{\delta_{x}}f=f(x)$ for all functions $f$.
@@ 3.1 Measure Theoretic Technicalities
The language as developed is fine for dealing with finite sets, where elements of the dual are merely finite dimension vectors. When dealing with infinite sets, the issue of how to represent elements of $\left(\mathbb{R}^{X}\right)^{*}$ becomes more troublesome. The usual route is to only consider loss profiles that are measurable. Rather than all functions $\mathbb{R}^{X}$ we work with a subalgebra of measurable functions, which we call $\mathrm{Meas}(X,\mathbb{R})$. A probability distribution is then a positive, normalized, linear function $\mathrm{Meas}(X,\mathbb{R})\rightarrow\mathbb{R}$. Restricting oneself in this way affords the representation of probability distributions via the Lebesgue integral. Of course, one could take our work as pertaining to $\mathrm{Meas}(X,\mathbb{R})$ rather than all of $\mathbb{R}^{X}$, without any real loss in the results. Our primary interest is when:
The Basic Sets of Study are Finite.
@@
At times, the natural set of actions will not be a finite set, but rather a well-behaved subset of $\mathbb{R}^{n}$.
@@ 4 Optimal Decisions
Returning to our decision maker, we assume that they represent their uncertainty in $\Theta$, via a probability distribution $P\in\mathbb{P}(\Theta)$, and that they choose their actions minimizing the expected loss.
@@ 4.1 Representing Loss Functions
Here we develop general representations of loss functions $L:\Theta\times A\rightarrow\mathbb{R}$. We assume that the sets $\Theta$ and $A$ are finite. Even if one starts with a finite set of actions, in many statistical problems it is natural to consider distributions over unknowns $\mathbb{P}(\Theta)$.
@@
A proper loss takes a prediction $Q\in\mathbb{P}(\Theta)$, and then penalizes the decision maker according to the amount of weight that their prediction assigns to the unknown $\theta$. Properness ensures that if the decision maker knows $P$, then they minimize their expected loss by reporting $P$. Proper losses constitute a well-studied class of loss functions that provide suitable surrogates for decision problems [5, 10, 22, 9, 6, 15, 6, 1].
Furthermore, we show how to render any proper loss convex through a canonical reparameterization. This allows the use of tools from convex analysis [4, 14] to aid in calculating optimal actions.
@@ 4.2 Entropy from Loss
Rather than working with probability distributions, we take the route of [21] and work with unnormalized distributions. Denote the set of all unnormalized distributions on $\Theta$ by $\mathbb{P}^{+}(\Theta)$. For any loss function $L$, define the entropy ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$,
@@
A function $f:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ is 1-homogeneous if for all $\mu\in\mathbb{P}^{+}(\Theta)$ and for all $\lambda>0$,
$f(\lambda\mu)=\lambda f(\mu).$
@@ 4.3 Loss from Entropy
All loss functions give rise to an entropy. The entropy encodes much information about its associated loss through its super-gradients, which includes all the Bayes actions for the underlying loss. For any distribution $P$, define the Bayes actions for $P$ as the set of minimizers,
@@
Denote the set of all supergradients at a point $x$ by $\partial f(x)$, and the set of all supergradients by $\partial f=\cup_{x}\partial f(x)$. For differentiable concave functions, the supergradients are the same as regular gradients [14]. 1-homogeneous functions afford a very simple representation via their super-gradients.
@@
Let $f:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave 1-homogeneous function. Then for all $x$ and for all $v\in\partial f(x)$,
@@
which follows directly from the definition of a supergradient at $x$ and the 1-homogeneity of $f$. Rearranging yields $\frac{1}{2}(f(x)-\left<x,v\right>)\geq 0$, which implies $f(x)\geq\left<x,v\right$>. Similarly,
@@
which follows directly from the definition of a supergradient at $\lambda x$ and the 1-homogeneity of $f$. Rearranging yields $\frac{1}{2}(f(x)-\left<x,v\right>)\leq 0$, yielding $f(x)\leq\left<x,v\right$>. Combing the results gives $f(x)=\left<x,v\right$>. To prove the second claim, we have for all $y$ and $\lambda>0$,
@@
where the first line is by definition and the second is by 1-homogeneity. As $y$ is arbitrary, the claim is proved.
This theorem provides a corollary that shows the supergradients of a 1-homogeneous function have a property similar to properness.
@@
We now show that the partial loss of a Bayes action is a supergradient of $\underline{L}$}.
@@
Let ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave 1-homogeneous function. Then its canonical loss, ${\mathcal{L}}:\Theta\times\partial{\underline{L}}\rightarrow\mathbb{R}$ is given by ${\mathcal{L}}(\theta,\zeta)=\zeta(\theta)$.
As will be shown, canonical losses can always be convexified. Furthermore, they maintain all the properties of $L$ needed to assess the quality of decisions.
4.3.1 The Bayes Super-Prediction Set
The process of canonising a loss, that is, going from
@@ 4.3.1 The Bayes Super-Prediction Set
can create extra partial losses/actions that were not originally available to the decision maker. However, they do not get benefit from these extra actions. From any entropy define the Bayes super prediction set,
@@
The Bayes super-prediction set is precisely those partial losses that the decision maker does not need to use over the actions available to them, regardless of the distribution $P$. The super-prediction set is convex. Furthermore, the Bayes actions for $\mathcal{L}$} are the lower boundary of the super prediction set.
@@
Let ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave 1-homogeneous function. Then $\zeta\in\partial{\underline{L}}$ if and only if,
@@
The proof is a straightforward application of 1-homogeneity and supergradients.
@@
Let ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave 1-homogeneous function and let ${\nabla{\underline{L}}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}^{\Theta}$ be a super-gradient function, ${\nabla{\underline{L}}}(\mu)\in\partial{\underline{L}}(\mu),\ \forall\mu$. Then,
@@
is a proper loss.
@@ 4.4 Convexification of Losses in Canonical Form
The preceding shows how to construct losses, we begin with a concave 1-homogeneous function and take super-gradients. The focus now turns to their convexification. Once convexified, the decision maker has access to the large and ever-growing literature on the minimization of convex functions to aid in the calculation of optimal actions. We closely follow [6], with a focus on canonical losses. This streamlines the development. For example, for some proper losses, lemma 12 fails to hold. Furthermore, our result on convexification of canonical losses (theorem 14), is to the best of our knowledge novel.
Recall $\bm{1}_{\Theta}\in\mathbb{R}^{\Theta}$ is the function that always returns $1$, and define $\bm{1}_{\Theta}^{\bot}$ to be its orthogonal complement in $\mathbb{R}^{\Theta}$, the functions $v\in\mathbb{R}^{\Theta}$ with,
$\left<\bm{1}_{\Theta},v\right>=\sum_{\theta\in\Theta}v(\theta)=0.$
@@
Proof Fix $v$ and suppose that there exist $\gamma_{1}$ and $\gamma_{2}$ with $\gamma_{1}<\gamma_{2}$ and $\gamma_{1}\bm{1}_{\Theta}+v,\gamma_{2}\bm{1}_{\Theta}+v\in\partial{\underline{L}}$. By assumption, $\gamma_{2}\bm{1}_{\Theta}+v$ is Bayes for some distribution $P$. But,
@@
Thus, we lose nothing by working with projections of losses onto $\bm{1}_{\Theta}^{\bot}$. Define,
@@
Proof To show that $\hat{\Gamma}_{{\underline{L}}}$ is convex, we are required to show that for all $\zeta_{1},\zeta_{2}\in\partial{\underline{L}}$ and all $\lambda\in[0,1]$ there is a constant $\gamma$ such that,
@@
with equality for $P^{*}$. Therefore, by lemma 9, $\lambda\zeta_{1}+(1-\lambda)\zeta_{2}-\gamma^{*}\bm{1}_{\Theta}\in\partial{\underline{L}}$.
@@
Let ${\underline{L}}:\mathbb{P}^{+}(\Theta)\rightarrow\mathbb{R}$ be a concave 1-homogeneous function. Then its canonical loss $\mathcal{L}$} can be represented as ${\mathcal{L}}:\Theta\times C\rightarrow\mathbb{R}$, with $C\subseteq\bm{1}_{\Theta}^{\bot}$ a convex set and,
${\mathcal{L}}(\theta,v)=\left<\delta_{\theta},v\right>+\Psi(v),$